If you start with 16 million nuclei of a radioisotope with a half life of 10 minutes, how many will be left after 110 minutes?
Answer #1

N = Noe-\lambdat

T1/2 = ln 2/\lambda

\lambda = ln 2/T1/2 = ln 2 /10 = 0.0693147

N = 16*e-0.0693147*110 = 0.0078125 millions